There are tree triplets. Ed and Ned have $135 total combined, Ed and Ted have $153 total combined, and Ned and Ted have $122 combined. How much do all three have together? And how much do each individual have?
Answer by ankor@dixie-net.com(22740)

There are tree triplets. Ed and Ned have $135 total combined, Ed and Ted have $153 total combined, and Ned and Ted have $122 combined. How much do all three have together? And how much do each individual have?
:
Using elimination, write the three equation and add as follows
e + n + 0 = 135
e + 0 + t = 153
0 + n + t = 122
-----------------
2e + 2n + 2t = 410
Simplify, divide by 2
e + n + t = 205
:
Subtract the 1st equation from the above
e + n + t = 205
e + n + 0 = 135
-------------------Subtaction eliminates e and n, find t
0 + 0 + t = 70
t = $70 is Teds amt
:
Using the 3rd equation, replace t with 70
n + 70 = 122
n = 122 - 70
n = $52 is Ned's amt
:
Used 2nd equation, replace t with 70 again
e + 70 = 153
e = 153 - 70
e = $83 is Ed's amt
:
:
You can confirm this by adding these value together and see if they agree with the given sums